OpenStudy (anonymous):

Solve the equation. Express numbers as integers or simplified fractions. Separate multiple answers with a comma, if necessary. 0=4t^3-5t^2-36t+45

OpenStudy (anonymous):

\[t^2\left( 4t-5 \right)-9\left( 4t-5 \right)=0\] \[\left( 4t-5 \right)\left( t^2-9 \right)=0\] \[\left( 4t-5 \right)\left( t+3 \right)\left( t-3 \right)=0\] find x